OpenStudy (anonymous):

1/4 (2x − 14) = 4 ? show me how you got the asnwer too please not just the answer

OpenStudy (anonymous):

Distribute 1/4 to the equation. Giving .5x - 3.5 = 4 .5x - 3.5 + 3.5 = 4 + 3.5 .5x/.5 = 7.5/.5 x = 15

OpenStudy (anonymous):

5/2x + 15/2 = 15 ?

OpenStudy (anonymous):

thans so much your ahuge help :)